OpenStudy (anonymous):

Ken has only two girl cousins, Michelle and Laura, and the rest are boys. The sum of Michelle and Laura's ages is 22. Laura is 2 years older than Michelle. How old is Michelle?

OpenStudy (anonymous):

i need help

OpenStudy (aripotta):

M + l = 22 L = m + 2

OpenStudy (anonymous):

the answer is?

OpenStudy (aripotta):

Substitute m + 2 for l in the first equation and solve for m

OpenStudy (anonymous):

thanks is

OpenStudy (aripotta):

So m + (m + 2) = 22

OpenStudy (anonymous):

x+(X+2)=22 2x+2=22 2x=20 x=10 one cousin is 10, other x+2=12 yrs. old